OpenStudy (anonymous):

I'd like to prove this infinite series formula. Any hints will be very appreciated, thanks. (Writing the sum in the first post)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}nr^n=\frac{ r }{ (1-r)^2 }\]

terenzreignz (terenzreignz):

Are you sure it's r and not nr on the numerator?

terenzreignz (terenzreignz):

ok, sorry, it is r, after all. It's all about sorting them, I guess?

terenzreignz (terenzreignz):

You know the formula for the sum of an infinite geometric series?

terenzreignz (terenzreignz):

Ok, so our series takes the form \[r^{1} + 2r^{2} + 3r^{3} + 4r^{4}...\] this can be written as \[r^{1} + r^{2} + r^{3} + r^{4}...\]\[+\]\[r^{2}+r^{3}+r^{4} + r^{5}...\]\[+\]\[r^{3}+r^{4}+r^{5} + r^{6}...\]\[+\]\[r^{4}+r^{5}+r^{6} + r^{7}...\]\[...\] Each of these is just a geometric series with common ratio r, they only differ in their first terms

terenzreignz (terenzreignz):

Them being geometric series, their sums are given by... \[\frac{r}{1-r}+\frac{r^{2}}{1-r}+\frac{r^{3}}{1-r}+\frac{r^{4}}{1-r}...\] Which is just \[\sum_{n=1}^{\infty}\frac{r^{n}}{1-r}\] voila, another geometric series :)

OpenStudy (anonymous):

Ah, amazing, thanks!

terenzreignz (terenzreignz):

No problem :D

OpenStudy (anonymous):

The rest being, of course: \[\sum_{n=1}^{\infty}=\frac{ r^n }{ 1-r }=\frac{ 1 }{ 1-r }\sum_{n=1}^{\infty} r^n=\frac{ 1 }{ 1-r }\frac{ r }{ 1-r }=\frac{ r }{ (1-r)^2 }\]

terenzreignz (terenzreignz):

That's right :D